State whether or not the following functions have a well-defined inverse. If the inverse is well-defined, define it. If it is not well-defined, provide justification.

a)    f : Z → Z. f(x) = 7x – 7

b)    f : R → R. f(x) = 7x – 7

c)    A = {a, b, c, d, e}. f : P (A) → {0, 1, 2, 3, 4, 5}. f(x) = |x|. It maps a set to the number of elements it contains.


1
Expert's answer
2021-11-22T15:06:08-0500

a) "f: z \rightarrow z \quad f(x)=7 x-7"

Domain is integer, codomain is integer

Take "\quad x=1, \quad \Rightarrow f(x)=0"

 "x=2, \quad \Rightarrow \quad f(x)=7"

Thus, "\quad f(x) \neq 1,2,3,4,5 \ldots"

Range is not same as codomain. It is not onto. Thus, no inverse.

b) "f: R \rightarrow R \quad f(x)=7 x-7"

For "R \rightarrow R \quad f(x)" is both one-one and onto.

Thus, inverse exist.

"y=7 x-7\\\Rightarrow x=\frac{y+7}{7}\\ \therefore \quad f^{-1}(x)=\frac{x+7}{7}"

c) |x| is not a one-one function. It has no inverse.